5. Suppose that X X, and X are independent random variables such that ELX.]-O and BLX?) -1 for is 12.3. Find the value of EIX/(2x1-X3M
Answer #1

from above E(X22*(2X1-X3)2)=E(X22*(4X12+X32-4X1X3))

=4*E(X22)*E(X1)2+E(X22)*E(X3)2-4E(X22)*E(X1)*E(X3)

=4*1*1+1*1-4*1*0*0=5